UT Feb 26 2021



Night Partition

  • Night split with LBTI - night center 00:32 - handover 00:27 to LBTI
  • Moon is up effectively all night (98.4%)
  • 1st half D/E-time
    • Twilight (19:15hr)-22:15 - HD 33632 - LBTI w/ALES
    • MODS Field Correction check for IT 8315 - 22:30-00:27 - MODS
      • MODS or LBC science could replace this slot
  • 2nd half LBTI
NO LUCI-1 SOUL AO

Night Plan for D/E time

  • Engineering Tasks
    • MODS Field Correction check for IT 8315
      • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
      • priority=medium (problem may have been with us for a year)
      • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
        • do LDG and RDG in parallel (two IDL windows)
        • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
        • Stone_G is well placed at RA=10:17, and other Stone / BS fields are available if earlier or later.
  • Science
    • HD33632 brown dwarf companion observation - LBTI/ALES - (Stone) - 3hr
      • 05 13 17 +37 20 14
      • Directors time from December, lost to weather and M3 issue
      • Target from twilight + 3hrs
      • Weather: clear, <1.2-ish arcsec seeing
    • Nova - 0.5hr MODS-Bino (Wagner)
      • Any Conditions
    • Local Merger min 0.5hr MODS-Bino (Rothberg)
      • Any Conditions
      • Can be used as a filler if needed
      • at least 1x1800sec exposure in all 4 MODS channel - but red is the highest science priority (in case something breaks...)
      • NGC 3310 - large target 36deg from Moon - Red channel is highest priority
      • script location: Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts

Overview

20210225 23:21 Starting up the LBTI WFS, MODS, and all of the other toys we will be using tonight. Both MODS awake and simSnap run

20210226 0:37 David has completed his checkouts and test preset. Running some mods calibs while we wait for sunset.
  • MODS1 & 2 arcs mods{1|2}r/b.20210226.0003-0005
  • MODS1 & 2 slitless flats mods{1|2}b.20210226.0006-0015, mods{1|2}r.20210226.0006-0010
20210226 1:17 Sunset

20210226 1:20 David is opening

20210226 1:30 Slewing to Alpha Aur for setup
  • David is working on setting the correct flag for the syncronous preset
  • Aligned M3 with cold stop Tip 45" Tilt 15"
  • and setup AO, aligned sweetspots for ALES. Ilya set up PEPSI. Even the few minutes taken to set up got a very high s/n spectra!
2:00 Slewing to HD33632. See LBTI nightlog for details.
https://docs.google.com/spreadsheets/d/1c1RKNSMQWG1NowLX9f96MOEzyrWUEj8974ao9IfdhQ4/edit?usp=sharing

Seeing is not great ~1.3-5" on the DIMM, 1-1.3" on PEPSI guider. And there appears to be periodic "invisible cloud cover" wafting through (short blimps of decreased counts seen on WFS that correspond to water vapor spikes in ALES spectrum). All sky camera is perfectly clear. Conditions still look ok to move forward with observation.

20210226 2:07 12 deg Twilight

Seeing highly variable 1.2-2" with some water vapor despite 4% humidity at LBT. SMT water vapor is all over the place so somethign there. We are still gettign what looks to be useful data.

20210226 3:00 DX M1 has panic'd we were going to continue to track but the seeing has blown right up to 3.5" seeing. This data is not useful at all. David is taking us up to recover. Steve is relinquishing the telescope for other tasks since this is no longer even emotely useful. David is reconfiguring for MODS/MODS after he recovers the DX primary.

20210226 3:30 Pointing check. Seeing is 4.3"....actually 5". Very abrupt temperature drop from +1.3 to -3.3C and continuing to drop. The star on GCS is larger than the guide FoV.
Conditions are non photometric. David has verified what we suspected suring the LBTI obs, there are some thing spots of high thin cirrus. So non photometric. Doing phtometric standards does not make sense. We can not guide on this, can not make out a 13th mag star on the guider. We are in weather downtime until we can make something out.

20210226 4:15 Sending Preset to G191B2B seeing was <4", and I could make out the GS somewhat. We will give this a shot.

Seeing 3.4-4.0" mods{1|2}b.20210226.0016-0018, mods1r.20210226.0013-0015, mods2r.20210226.0015-0017

20210226 4:38 Going to V1112Per. We will see.... and the seeing is 5" again. The guide stars landed in the patrol field but then blew up to indistinguishable. Holding for a moment.

20210226 4:49 Sent again. Seeing seems to be "settling down" to about 3". These are certainly not the intended conditions for this, but grabbing V1112 science. Getting some extra exposures. Seeing is bouncing betwen 2.4-3.5" on the DIMMfor first exp. Very strong and obvious emission lines so we are good I think.
  • PA set to 85. Slit to 1.2" NEXP6
  • 5:20 Seeing improving 1.5-2.5", better towards the last exposures.
  • mods{1|2}b.20210226.0019-0024, mods1r.20210226.0019-0024, mods2r.20210226.0023-0028
20210226 5:42 Preset to NGC3310 for Local Mergers DG science.
  • Seeing 1.7-2.8"
  • 6:00 Seeing is really going to pot again.... temperature has risen back up to 1.1C again very suddenly. Still aligning MODS2 when things blew up... but MODS1 integrating.
  • 6:11 OK it looks like it might be stabilizing back to 2.8-3"...MODS 2 is just a little behind (my bad, thought we might be stopping for poor weather again). We can take some additional time on MODS1 at the end to even that out.
  • 6:30 Seeing 3.0-4.2". My vote is stay and exposure here on this until we handover and hope we get something... This is really weather downtime
  • 6:48 Seeing 2.47-2.85". Steve is on for LBTI. We will stay here until reconfig time.
  • mods{1|2}b.20210226.0025-0026, mods1r.20210226.0028-0029, mods2r.20210226.0033-0034
  • partials mods{1|2}b.20210226.0027, mods1r.20210226.0030 mods2r.20210226.0035
20210226 7:24 preparing for a reconfig from MODS MODS to LBTI/PEPSI. Kevin, Ilya, Jared and Steve are online for the LBTI time

20210226 7:32 Night Center

20210226 8:15 Seeing is improving for LBTI! Lucky. 1.3-2.0". They are set up and working on AZ program for Kevin Wagner. MODS has been put to sleep. Pending calibrations for MODS for programs executed will be completed tomorrow.

20210226 13:47 Sunrise

-- BarryRothberg - 25 Feb 2021
Topic revision: r4 - 26 Feb 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback